# Large Language Models for Crash Detection in Video: A Survey of Methods, Datasets, and Challenges

### v1

@@ II-C Large Language Models for Video Understanding
II-C1 Multimodal Extensions of LLMs
@@
II-C2 Applications in Crash Detection
@@
II-C3 Advantages and Integration with VLMs
LLMs offer distinct advantages over classical methods and VLMs by providing contextual understanding and explainability. While classical approaches struggle with ambiguous scenarios [6], and VLMs like Video-LLaMA focus on spatiotemporal feature alignment [27], LLMs excel in reasoning about high-level semantics and causality. For example, integrating LLMs with VLMs, as in VERA, combines LLaVA-1.5’s visual processing with verbalized reasoning to detect and explain crash events (e.g., “a car collided due to sudden braking”) [29]. This synergy enables robust real-time crash detection, addressing challenges like varying lighting or complex traffic interactions [10]. Future research should focus on optimizing LLM adapters for real-time processing and developing comprehensive video crash datasets to further enhance performance.
II-D Taxonomy of LLM-Based Crash Detection
The integration of Large Language Models (LLMs) into crash detection systems has led to diverse approaches, categorized along three primary axes: fusion level, prompt strategy, and LLM role. Additionally, we distinguish systems based on input granularity (frame-level vs. clip-level encoding) and detection focus (causal detection vs. post-event summarization). This taxonomy provides a structured framework to analyze LLM-based crash detection methods, highlighting their technical nuances and applicability to real-world traffic scenarios [11, 33]. By building on classical approaches [3] and Vision-Language Models (VLMs) [27], LLM-based systems leverage multimodal reasoning to address the variability and ambiguity of crash events.
II-D1 Fusion Level
@@
Vision Encoder Visual Tokens Text Encoder Text Tokens Concatenate LLM (a) Early Fusion (e.g., Video-LLaMA) Vision Encoder Visual Features Text Encoder Text Features Fusion Module Output (b) Late Fusion (e.g., CrashLLM) Vision Encoder Visual Features LLM (c) Cross-Attention (e.g., VAD-LLaMA) Cross-Attention Layers
@@ II-D3 LLM Role
II-D4 Input Granularity and Detection Focus
@@
II-D5 Discussion
The taxonomy highlights the diversity of LLM-based crash detection approaches, each addressing specific challenges in video understanding. Early fusion and cross-attention strategies enhance multimodal integration, while dynamic prompts and generative roles improve contextual reasoning and explainability. Clip-level encoding and causal detection are particularly suited for real-time applications, overcoming limitations of classical methods like frame differencing [6, 10]. However, challenges such as computational complexity and dataset limitations persist [11]. Future research should focus on optimizing fusion strategies and developing comprehensive video crash datasets to advance LLM-based systems.
TABLE I: Comparison of Key Datasets for Video-Based Crash Detection and Analysis
Dataset Year Size Source Annotation Type Focus DAD [38] 2016 1,500+ videos Dashcam Frame-level crash events Supervised, multi-class crash detection (e.g., rear-end, side-impact). CADP [39] 2019 ~2,000 videos Police reports Crash causes, outcomes Temporal modeling and causal relationship analysis in crashes. Berkeley DeepDrive (BDD100k) [40] 2020 10,000+ clips Dashcam Weak labels (crash/anomaly) Anomaly detection and semi-supervised learning for autonomous driving. UCF-Crime [41] 2018 1,900 videos Surveillance Anomaly categories (subset) Benchmarking anomaly detection, including ”road accident” category. SHRP 2 NDS [42] 2024 (used) 8,600+ SCEs Naturalistic Driving Event/Conflict types Differentiating event severity (crash, near-crash) and conflict types. WTS [43] 2024 (used) 810+ videos Mixed (Overhead, Ego) Fine-grained captions Dense video captioning and multi-phase event description. CCD [44] 2024 (used) N/A Dashcam Crash events Crash recognition and anticipation from ego-vehicle perspective.
Size (log-scaled)Annotation GranularitySource Diversity1111222233334444DADCADPBDD100kUCF-CrimeSHRP 2 NDSWTS
Figure 4: Radar chart comparing key attributes of crash detection datasets. ’Size’ is shown on a normalized log scale, ’Annotation Granularity’ is on a scale from 1 (weak) to 4 (fine-grained), and ’Source Diversity’ is on a scale of 1 (single) to 2 (mixed).
@@ III Datasets and Benchmarks
The development of Large Language Model (LLM) and Vision-Language Model (VLM)-based crash detection systems relies on high-quality datasets that provide annotated video data, temporal sequences, and contextual information for training and testing. These datasets, summarized in Table I, enable models to identify crash events in diverse traffic scenarios, addressing limitations of classical approaches [3]. Early datasets like DAD and UCF-Crime provided foundational benchmarks, while recent additions like SHRP 2 NDS [42] offer much larger scale and finer-grained annotations, crucial for training robust multimodal models [11, 42]. Additionally, newer LLM-augmented datasets are emerging, incorporating synthetic captions, temporal questions, or causal chains to enhance multimodal reasoning and support advanced crash detection tasks [33, 27].
@@ III-B LLM-Augmented Datasets
The reviewed datasets provide a foundation for advancing LLM and VLM-based crash detection, offering diverse data from annotated dashcam videos (DAD) to large-scale naturalistic driving studies (SHRP 2 NDS). LLM-augmented datasets, with synthetic captions and causal chains, further enhance multimodal reasoning, addressing challenges like coarse annotations in UCF-Crime [41]. However, issues such as dataset imbalance, limited real-world diversity, and computational demands for processing large video datasets persist [11]. Future efforts should focus on developing comprehensive datasets with fine-grained annotations and expanding LLM-augmented datasets to support real-time crash detection.
@@ IV-A Visual Encoder + LLM Decoder
In this architecture, video frames are encoded using a vision encoder, such as Vision Transformer (ViT) or Inflated 3D ConvNet (I3D), and the resulting features are fed as tokens to an LLM decoder for multimodal reasoning. Inspired by models like Flamingo, this approach processes video frames as a sequence of visual tokens, enabling the LLM to generate crash-related predictions or descriptions [26]. For instance, Video-LLaMA employs a ViT-based encoder to extract frame features, which are tokenized and processed by an LLM to detect crash anomalies on datasets like UCF-Crime [27, 41]. This architecture excels in tasks requiring temporal reasoning, such as identifying crash sequences, but demands significant computational resources due to end-to-end processing of visual and textual data [11].
@@ IV-B Frozen LLM + Learned Adapter
This architecture keeps the LLM backbone, such as GPT-4 or LLaMA-2, frozen and trains lightweight adapters to bridge vision features and the language model. Adapters, as used in BLIP-2 and MiniGPT-4, enable efficient fine-tuning for crash detection without modifying the pre-trained LLM [25, 45]. For example, CrashLLM employs a LoRA adapter to integrate video-derived features from a Swin Transformer with a frozen LLaMA-2 backbone, achieving high accuracy on the CrashEvent dataset [33, 32, 31]. This approach is computationally efficient, making it suitable for real-time applications, and supports tasks like crash outcome prediction and causal inference [35].
@@ IV-C Joint Vision-Language Pretraining
Joint vision-language pretraining involves training both vision and language components together on crash-like tasks, enabling end-to-end optimization for crash detection. Due to data scarcity, few public examples exist, but frameworks like VERA demonstrate this approach by pretraining on crash-specific datasets like DAD and CADP [29, 38, 39]. These models combine vision encoders (e.g., SlowFast) and LLMs (e.g., GPT-3.5) to learn crash-specific patterns, such as sudden vehicle movements or collisions [46, 30]. Joint pretraining enhances robustness in diverse traffic scenarios but requires large, annotated datasets, limiting its adoption [11]. Efforts to address data scarcity include synthetic data generation [33]. A recent trend involves hybrid approaches, such as ScVLM [42], which combines supervised learning for event classification with contrastive learning for nuanced event and conflict type identification, leveraging the strengths of both paradigms.
@@ IV-D3 Proposed Framework and Evaluation
A powerful approach would be to combine SSL and FSL. A backbone model like LLaMA-2 [31] could be pre-trained using SSL on a large corpus of unlabeled dashcam videos. This pre-trained model would develop a rich understanding of general traffic scenes. Subsequently, it could be fine-tuned for specific crash detection tasks using an FSL approach with just a few labeled examples from a dataset like DAD [38]. Such a model, which we refer to as SSL-VAD in Table II, would be highly data-efficient and adaptable to new, rare crash scenarios. Its performance could be evaluated on standard benchmarks like UCF-Crime, with expected performance rivaling supervised methods due to the powerful representations learned during SSL pre-training.
@@ IV-E Common Components
The three architectures—Visual Encoder + LLM Decoder, Frozen LLM + Learned Adapter, and Joint Vision-Language Pretraining—offer distinct trade-offs for crash detection. The Visual Encoder + LLM Decoder excels in temporal reasoning but is computationally intensive, while Frozen LLM + Learned Adapter provides efficiency for real-time systems. Joint Vision-Language Pretraining promises robustness but is limited by data scarcity [11]. The enhanced common components—ViT, Swin, SlowFast for vision; Transformer, ConvLSTM, TimeSformer for temporal modeling; and GPT-3.5/4, LLaMA-2, MiniGPT-4 for language—enable these architectures to surpass classical methods like frame differencing in handling complex crash scenarios [6, 10]. Future research should focus on optimizing component efficiency, developing hybrid architectures, and addressing data scarcity through synthetic datasets to advance real-time crash detection.
@@ V-A Crash Detection Accuracy
Crash Detection Accuracy measures a model’s ability to correctly classify video segments as containing a crash event, either as a binary (crash vs. non-crash) or multi-class (e.g., rear-end collision, side impact, pedestrian accident) classification task. It is typically reported as the percentage of correct predictions over a test set. For instance, VERA achieves 86.55% accuracy on the UCF-Crime dataset for binary crash detection, while CrashLLM reports over 92% accuracy on multi-class classification using the CrashEvent dataset [29, 33, 41]. This metric is critical for evaluating model robustness across diverse crash scenarios, particularly when trained on datasets like DAD or CADP [38, 39]. However, accuracy alone may be insufficient for imbalanced datasets, where metrics like F1-score, which considers both precision and recall, are often used to account for false positives and negatives [11]. The Area Under the ROC Curve (AUC) is another widely used metric that evaluates the model’s ability to distinguish between classes across all classification thresholds [9].

### v2

@@ II-D Large Language Models for Video Understanding
II-D1 Multimodal Extensions of LLMs
@@
II-D2 Applications in Crash Detection
@@
II-D3 Advantages and Integration with VLMs
LLMs offer distinct advantages over classical methods and VLMs by providing contextual understanding and explainability [57]. While classical approaches struggle with ambiguous scenarios [58], and VLMs like Video-LLaMA focus on spatiotemporal feature alignment [38], LLMs excel in reasoning about high-level semantics and causality. For example, integrating LLMs with VLMs, as in VERA, combines LLaVA-1.5’s visual processing with verbalized reasoning to detect and explain crash events (e.g., “a car collided due to sudden braking”) [54]. This synergy enables robust real-time crash detection, addressing challenges like varying lighting or complex traffic interactions. Future research should focus on optimizing LLM adapters for real-time processing and developing comprehensive video crash datasets to further enhance performance.
II-E Taxonomy of LLM-Based Crash Detection
The integration of Large Language Models (LLMs) into crash detection systems has led to diverse approaches, categorized along three primary axes: fusion level, prompt strategy, and LLM role. Additionally, we distinguish systems based on input granularity (frame-level vs. clip-level encoding) and detection focus (causal detection vs. post-event summarization). This taxonomy reveals a critical insight: the architectural choices made at each level fundamentally determine not just performance, but the system’s suitability for specific real-world applications. By building on classical approaches [49] and Vision-Language Models (VLMs) [38], LLM-based systems leverage multimodal reasoning to address the variability and ambiguity of crash events, but each design choice carries inherent trade-offs that practitioners must carefully consider.
II-E1 Fusion Level
@@
The choice of fusion level represents a fundamental trade-off between semantic depth and computational cost. Early fusion allows the LLM to access raw visual nuances but risks overwhelming it with irrelevant data, while late fusion is more efficient but may lose critical, low-level visual cues necessary for distinguishing near-misses from actual crashes.
@@ II-E3 LLM Role
The LLM’s role directly maps to the system’s intended application. A ‘passive’ captioning model is sufficient for offline data logging, whereas an ‘active’ QA model is necessary for forensic analysis. A ‘generative’ reasoning model, while most insightful, carries the highest risk of factual hallucination, making it a double-edged sword for legal and insurance purposes.
II-E4 Input Granularity and Detection Focus
@@
II-E5 Discussion
The taxonomy highlights a fundamental paradox in current LLM-based crash detection: the most sophisticated approaches (early fusion with generative reasoning) offer the richest understanding but are least practical for deployment, while the most deployable approaches (late fusion with passive roles) sacrifice the very contextual understanding that makes LLMs valuable. This creates a critical research imperative: developing architectures that can bridge this gap between sophistication and practicality. Early fusion and cross-attention strategies enhance multimodal integration, while dynamic prompts and generative roles improve contextual reasoning and explainability. However, the computational complexity and dataset limitations reveal that the field is still in its early stages, with significant engineering challenges ahead [51].
TABLE II: Comparison of Key Datasets for Video-Based Crash Detection and Analysis
Dataset Year Size Source Annotation Type Focus DAD [61] 2016 1,500+ videos Dashcam Frame-level crash events Supervised, multi-class crash detection (e.g., rear-end, side-impact). CADP [62] 2019 ~2,000 videos Police reports Crash causes, outcomes Temporal modeling and causal relationship analysis in crashes. Berkeley DeepDrive (BDD100k) [63] 2020 10,000+ clips Dashcam Weak labels (crash/anomaly) Anomaly detection and semi-supervised learning for autonomous driving. UCF-Crime [64] 2018 1,900 videos Surveillance Anomaly categories (subset) Benchmarking anomaly detection, including “road accident” category. SHRP 2 NDS [65] 2024 (used) 8,600+ SCEs Naturalistic Driving Event/Conflict types Differentiating event severity (crash, near-crash) and conflict types. WTS [66] 2024 (used) 810+ videos Mixed (Overhead, Ego) Fine-grained captions Dense video captioning and multi-phase event description. CCD [67] 2024 (used) N/A Dashcam Crash events Crash recognition and anticipation from ego-vehicle perspective.
Size (log-scaled)Annotation GranularitySource Diversity11223344DADCADPBDD100kUCF-CrimeSHRP 2 NDSWTS
Figure 5: Radar chart comparing key attributes of crash detection datasets. ‘Size’ is shown on a normalized log scale, ‘Annotation Granularity’ is on a scale from 1 (weak) to 4 (fine-grained), and ‘Source Diversity’ is on a scale of 1 (single) to 2 (mixed).
@@ III Datasets and Benchmarks
The development of Large Language Model (LLM) and Vision-Language Model (VLM)-based crash detection systems relies on high-quality datasets that provide annotated video data, temporal sequences, and contextual information for training and testing. These datasets, summarized in Table II, enable models to identify crash events in diverse traffic scenarios, addressing limitations of classical approaches [49]. Early datasets like DAD and UCF-Crime provided foundational benchmarks, while recent additions like SHRP 2 NDS [65] offer much larger scale and finer-grained annotations, crucial for training robust multimodal models [51, 65]. Additionally, newer LLM-augmented datasets are emerging, incorporating synthetic captions, temporal questions, or causal chains to enhance multimodal reasoning and support advanced crash detection tasks [53, 38].
@@ III-A Key Datasets
A holistic view of these key datasets reveals a critical gap: a lack of ‘causal’ annotations at scale. While datasets like DAD provide frame-level crash events and CADP offers post-hoc police report causes, there is no large-scale dataset that explicitly links fine-grained temporal events (e.g., ‘driver looks at phone’) to the crash outcome within the video itself. This scarcity of annotated causal chains is the single biggest bottleneck preventing current models from moving beyond correlation to true causal reasoning, a prerequisite for effective crash anticipation systems.
@@ III-B LLM-Augmented Datasets
These datasets enable robust LLM-based crash detection, though challenges remain including dataset imbalance, limited diversity, and computational demands [51]. Future work should focus on comprehensive datasets with fine-grained annotations.
@@ IV Model Architectures
TABLE III: Technical Comparison of VLM Architectures for Crash Detection
Feature Visual Encoder + LLM Decoder Frozen LLM + Learned Adapter Joint Vision-Language Pretraining
Example Model(s) Video-LLaMA, Flamingo BLIP-2, MiniGPT-4, CrashLLM VERA, Unified-IO
Key Innovation End-to-end multimodal reasoning by treating video frames as a foreign language for the LLM. Parameter-Efficient Fine-Tuning (PEFT) via lightweight bridges. Deep, synergistic fusion of modalities learned from scratch.
Architectural Detail A vision encoder (e.g., ViT-G) output is projected to the LLM’s word embedding space. The full model is fine-tuned. A small network (e.g., Q-Former, LoRA) aligns frozen vision features with a frozen LLM. A unified Transformer encoder-decoder trained on interleaved image-text data.
Trainable Parameters Very High (e.g., >>8B for a ViT-G + 7B LLM). Very Low (e.g., 1M–100M for adapters; >>99% frozen). Extremely High (often >>10B).
[…]
Inference Speed Moderate to Slow. Depends on vision encoder and LLM size. Fast. Often uses smaller LLMs. Moderate to Slow. Often large, unified models.
Pros High performance ceiling due to deep, end-to-end optimization. Highly efficient to train; easily adapts pre-trained LLMs; lowers barrier to entry for research. Potentially the most robust and specialized for the target domain.
Cons Computationally expensive; prone to catastrophic forgetting; requires large annotated datasets. Performance may be capped by frozen components’ knowledge; modality misalignment risk. Requires massive, domain-specific datasets (often unavailable); highest computational cost.
@@ IV-A Visual Encoder + LLM Decoder
In this architecture, video frames are encoded using a vision encoder, such as Vision Transformer (ViT) or Inflated 3D ConvNet (I3D), and the resulting features are fed as tokens to an LLM decoder for multimodal reasoning. Inspired by models like Flamingo, this approach processes video frames as a sequence of visual tokens, enabling the LLM to generate crash-related predictions or descriptions [18]. For instance, Video-LLaMA employs a ViT-based encoder to extract frame features, which are tokenized and processed by an LLM to detect crash anomalies on datasets like UCF-Crime [38, 64]. This architecture excels in tasks requiring temporal reasoning, such as identifying crash sequences, but demands significant computational resources due to end-to-end processing of visual and textual data [51]. Basically, the core mechanism is a projection layer that transforms the high-dimensional visual embeddings into a sequence of tokens compatible with the LLM’s input dimension.
The vision encoder outputs a sequence of feature vectors, one for each patch or frame. A linear layer or a small multi-layer perceptron (MLP) then projects these vectors. The computational burden is immense; fine-tuning this entire pipeline involves backpropagating gradients through billions of parameters across both the vision and language models. This requires significant memory to store activations and gradients, often necessitating model parallelism techniques across multiple GPUs.
Application in Crash Detection: This architecture is best suited for academic or research settings where the primary goal is to push the boundaries of performance and narrative generation, with less emphasis on immediate deployment. Its high computational cost and need for massive datasets make it impractical for real-world, real-time systems today. However, this architecture’s strength lies in its ability to learn complex temporal relationships from scratch, making it suitable for tasks requiring detailed narrative generation of a crash sequence.
@@ IV-B Frozen LLM + Learned Adapter
This architecture keeps the LLM backbone, such as GPT-4 or LLaMA-2, frozen and trains lightweight adapters to bridge vision features and the language model. Adapters, as used in BLIP-2 and MiniGPT-4, enable efficient fine-tuning for crash detection without modifying the pre-trained LLM [19, 68]. For example, CrashLLM employs a LoRA adapter to integrate video-derived features from a Swin Transformer with a frozen LLaMA-2 backbone, demonstaring F1 score (53.8%) improvements on the CrashEvent dataset [53, 52, 48]. This approach is computationally efficient, making it suitable for real-time applications, and supports tasks like crash outcome prediction and causal inference [56]. This is a form of Parameter-Efficient Fine-Tuning (PEFT). The adapter’s design is the key innovation.
BLIP-2 and MiniGPT-4 use a Querying Transformer (Q-Former) [19], a small transformer-based module. The Q-Former takes visual features from the frozen encoder and uses a set of learnable query vectors to “ask questions” of the visual input, extracting the most salient information into a compact sequence of soft-prompt tokens.
CrashLLM utilizes Low-Rank Adaptation (LoRA) [52], which injects trainable low-rank matrices directly into the attention layers of the frozen LLM. Instead of training the original weight matrix $W$ (size $d\times d$), LoRA trains two smaller matrices $A$ (size $d\times r$) and $B$ (size $r\times d$), where the rank $r\ll d$. The update is represented as $W^{\prime}=W+BA$. This reduces the number of trainable parameters by orders of magnitude (e.g., from billions to a few million).
Application in Crash Detection: This is currently the most practical and promising architecture for real-world deployment. Its efficiency and modularity are key advantages for creating systems that need to be deployed on edge devices or rapidly adapted to specific local traffic conditions with minimal data and computational overhead. This approach is computationally efficient, enabling fine-tuning on a single GPU in hours rather than days, and its modularity allows for easily swapping different LLMs or vision backbones.
@@ IV-C Joint Vision-Language Pretraining
Joint vision-language pretraining involves training both vision and language components together on crash-like tasks, enabling end-to-end optimization for crash detection. Due to data scarcity, few public examples exist, but frameworks like VERA demonstrate this approach by pretraining on crash-specific datasets like DAD and CADP [54, 61, 62]. These models combine vision encoders (e.g., SlowFast) and LLMs (e.g., GPT-3.5) to learn crash-specific patterns, such as sudden vehicle movements or collisions [69, 17]. Joint pretraining enhances robustness in diverse traffic scenarios but requires large, annotated datasets, limiting its adoption [51]. Efforts to address data scarcity include synthetic data generation [53]. A recent trend involves hybrid approaches, such as ScVLM [65], which combines supervised learning for event classification with contrastive learning for nuanced event and conflict type identification, leveraging the strengths of both paradigms. The core idea is that co-training enables the model to learn fundamental cross-modal connections (e.g., the visual signature of “sudden braking” and its linguistic description) that are more robust than those learned via fine-tuning alone.
Application in Crash Detection: While theoretically the most robust, this approach is best viewed as a long-term research goal. Its adoption is severely hampered by the lack of large-scale, high-quality annotated crash datasets. Until such datasets are created (perhaps synthetically), this architecture will likely be confined to labs at major corporations or well-funded research institutions. The immense computational cost of pretraining from scratch further limits its adoption, and current research often relies on proxy tasks or synthetic data generation [53] to make this approach feasible.
@@ IV-D3 Proposed Framework and Evaluation
A powerful approach would be to combine SSL and FSL. A backbone model like LLaMA-2 [48] could be pre-trained using SSL on a large corpus of unlabeled dashcam videos. This pre-trained model would develop a rich understanding of general traffic scenes. Subsequently, it could be fine-tuned for specific crash detection tasks using an FSL approach with just a few labeled examples from a dataset like DAD [61]. Such a model would be highly data-efficient and adaptable to new, rare crash scenarios. Its performance could be evaluated on standard benchmarks like UCF-Crime, with expected performance rivaling supervised methods due to the powerful representations learned during SSL pre-training.
@@ IV-E Common Components
The three architectures—Visual Encoder + LLM Decoder, Frozen LLM + Learned Adapter, and Joint Vision-Language Pretraining—offer distinct trade-offs for crash detection. The Visual Encoder + LLM Decoder excels in temporal reasoning but is computationally intensive, while Frozen LLM + Learned Adapter provides efficiency for real-time systems. Joint Vision-Language Pretraining promises robustness but is limited by data scarcity [51]. The enhanced common components—ViT, Swin, SlowFast for vision; Transformer, ConvLSTM, TimeSformer for temporal modeling; and GPT-3.5/4, LLaMA-2, MiniGPT-4 for language—enable these architectures to surpass classical methods like frame differencing in handling complex crash scenarios [58]. Future research should focus on optimizing component efficiency, developing hybrid architectures, and addressing data scarcity through synthetic datasets to advance real-time crash detection.
@@ V-A Crash Detection Accuracy
Crash Detection Accuracy measures a model’s ability to correctly classify video segments as containing a crash event, either as a binary (crash vs. non-crash) or multi-class (e.g., rear-end collision, side impact, pedestrian accident) classification task. It is typically reported as the percentage of correct predictions over a test set.This metric is critical for evaluating model robustness across diverse crash scenarios, particularly when trained on datasets like DAD or CADP [61, 62]. However, accuracy alone may be insufficient for imbalanced datasets, where metrics like F1-score, which considers both precision and recall, are often used to account for false positives and negatives [51]. The Area Under the ROC Curve (AUC) is another widely used metric that evaluates the model’s ability to distinguish between classes across all classification thresholds [76]. For instance, VERA achieves an AUC of 86.55% on the UCF-Crime dataset for binary crash detection [54, 64], while CrashLLM demonstrates substantial improvements on the CrashEvent dataset, boosting the average F1-score from 34.9% to 53.8% across crash classification tasks [53].